Kolby’s Korndogs is looking at a new sausage system with an installed cost of $910,000. This cost will be depreciated straight-line to zero over the project’s seven-year life, at the end of which the sausage system can be scrapped for $105,000. The sausage system will save the firm $193,000 per year in pretax operating costs, and the system requires an initial investment in net working capital of $51,000.

  

If the tax rate is 30 percent and the discount rate is 6 percent, what is the NPV of this project? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  

NPV            $
Answer #1

Initial investment = 910,000 + 51,000 = 961,000

Annual depreciation = 910,000 / 7 = 130,000

Operating cash flow from year 1 to year 7 = (Savings - depreciation)(1 - tax) + depreciation

Operating cash flow from year 1 to year 7 = (193,000 - 130,000)(1 - 0.3) + 130,000

Operating cash flow from year 1 to year 7 = 44,100 + 130,000

Operating cash flow from year 1 to year 7 = 174,100

Non operating cash flow in year 7 = Market value + NWC - tax(market value - book value)

Non operating cash flow in year 7 = 105,000 + 51,000 - 0.3(105,000 - 0)

Non operating cash flow in year 7 = 105,000 + 51,000 - 31,500

Non operating cash flow in year 7 = 124,500

NPV = Annuity * [1 - 1 / (1 + r)n] / r + FV / (1 + r)n - initial investment

NPV = 174,100 * [1 - 1 / (1 + 0.06)7] / 0.06 + 124,500 / (1 + 0.06)7 - 961,000

NPV = 174,100 * [1 - 0.665057] / 0.06 + 82,799.61065 - 961,000

NPV = 174,100 * 5.582383 + 82,799.61065 - 961,000

NPV = $93,692.23